# FlashVault: Versatile In-NAND Self-Encryption with Zero Area Overhead

### v1

@@ I Introduction
Today, governments and private enterprises leverage personal data to deliver highly customized policies and services [79, 20, 98]. In doing so, they often collect sensitive information closely tied to individual privacy such as financial records and biometric data. While such data serves as a critical resource for enabling personalized services, it also becomes a prime target for attacks, with severe consequences if leaked. For instance, in 2013 and 2017, massive data breaches at two major U.S. corporations, Target (a retailer) and Equifax (a credit reporting agency), respectively, exposed sensitive personal information of millions of individuals [19, 35]. These incidents resulted in credit card fraud and identity theft, with financial damages reaching hundreds of billions of dollars. To mitigate such illegal use and data misuse, many countries including the United States, the European Union, Japan, and South Korea have enacted data protection laws that mandate encryption and appropriate technical safeguards when storing personal data on storage devices [106, 28, 24, 57]. In parallel, the industry has adopted standards like TCG Opal [118] and IEEE 1667 [41] to enhance storage security.
@@
In accordance with data protection laws, many countries require or recommend that public institutions and private enterprises handling personal information adopt national standard cryptographic algorithms [106, 28, 24, 57].
However, current commercial SSDs currently support only specific algorithms, e.g., AES, SHA, and RSA [94, 70, 34, 102].
As a result, whenever unsupported algorithms are required, encryption is offloaded to the software stack on the host CPU [107].
This software-based encryption incurs I/O latency and may increase system-wide response time [16].
Such delays not only degrade the user experience in public and enterprise services but also cause significant disruption in latency-sensitive applications, such as high-frequency trading systems [71] and decentralized exchanges [114].
[Limitation 2] Unencrypted data may be stored in NAND flash memory. Modern SSDs are equipped with data encryption and SED features [68, 117].
However, the activation of these features is typically determined by user configuration [27, 112].
@@
Even when encryption is enabled, residual traces of cryptographic keys may remain in NAND flash due to garbage collection (GC) and wear-leveling (WL) during the key revocation process [121, 3].
In addition, if a power loss occurs, plaintext data residing in the DRAM of the SSD may be flushed to NAND flash [2, 128].
[Limitation 3] NAND flash lacks architectural protections against software-based and physical attacks. Since commercial SSDs do not support trusted execution environments within their NAND flash chips, attackers can manipulate the flash translation layer (FTL) mapping tables to distort the data organization or analyze FTL logs to retrieve or tamper with residual data that have been logically deleted but remain physically present [47].
In addition, attackers can connect NAND readers to the flash chip to extract unencrypted data, or remnants of data left behind due to GC and WL [68].
@@
For instance, plaintext data transmitted between the SSD and host may be intercepted through bus probing [122, 123], and cryptographic keys may be inferred or stolen through power analysis [91, 90] or electromagnetic analysis [30, 97] during CPU-based encryption processes.
@@
This paper focuses on the architecture of FlashVault and its performance. While security functionality is a fundamental requirement driven by legal and industrial standards, performance is also a critical factor for the practical deployment of secure storage systems in real-world environments [58, 47, 17]. Moreover, the structural characteristics of FlashVault inherently mitigate security vulnerabilities observed in conventional systems (Section II-C).
@@ II-A NAND Flash Memory Technology
The memory cells in early-generation SSDs are based on floating gate transistors (FGTs). An FGT is a complementary metal-oxide-semiconductor (CMOS) technology capable of holding electrical charges in an isolated silicon conductive layer (Fig. 2-(a)) [127]. Early generation flash memory arrays have a 2D planar structure with FGTs placed side-by-side on the die (Fig. 2-(c)). However, as the process node has scaled down, FGTs encountered fabrication challenges, i.e., charge leakage caused by the thin conductive layer and intensified cell-to-cell interference [60, 12]. To alleviate this limitation, FGTs were gradually superseded by transistors utilizing charge trap technology. These newer transistors employ a non-conductive silicon layer as a control gate, highly effective at retaining electrical charges (Fig. 2-(b)) [127, 12]. Moreover, to meet demands for higher density, memory suppliers developed a 3D NAND structure (Fig. 2-(d)), featuring vertically stacked charge trap transistors alongside the peripheral circuit on the same integrated chip [87, 113, 15]. This 3D architecture significantly increased storage capacity by adding more layers achieving lower cost/bit. More recently, the industry has introduced 4D NAND flash products111In the industry, 4D V-NAND is also referred to as PUC (Periphery Under Cell) by SK hynix, CUA (CMOS Under Array) by Micron, and COP (Cell Over Periphery) by Samsung [43]., which adopt an advanced vertical integration scheme to improve bit density (Fig. 2-(e)). These offer a reduced chip area compared to 3D NAND flash by positioning the peripheral circuit underneath the memory array [51]. Fig. 3-(a) illustrates the area composition of 512 Gb-class 3D V-NAND chips from three major memory manufacturers, indicating that peripheral circuits account for approximately 2024% of the total chip area. By placing the peripheral circuits under the memory array, the 4D structure not only reduces the die footprint by a comparable amount (Fig. 3-(b)), but also leaves a portion of the bottom silicon layer unoccupied, since the relocated circuits are smaller in area than the memory array above. Motivated by this architectural characteristic, there have been efforts to leverage the resulting unused area by integrating auxiliary logic circuits, such as pattern matchers [18], data processing units [108], and neuron-inspired circuits [39]. This work leverages this area to embed reconfigurable cryptographic engines, enabling fast and secure near-data cryptographic processing.
@@ II-B Internal Architecture of SSDs
Fig. 4 illustrates the typical architecture of modern SSDs [115, 12, 32, 48]. An SSD consists of NAND flash packages, an SSD controller, and DRAM. Each NAND flash package contains multiple flash chips, each housing several memory planes. These planes are further divided into blocks and pages. A block is the smallest erasable unit, typically holding 1,024 pages. A page, the basic unit for read and program/write operations, usually has a size between 4 KB and 16 KB. Data read from or programmed to the flash memory is managed by a flash memory controller. This controller orchestrates data distribution to prevent bottlenecks among NAND packages sharing a byte-wide channel. It also performs error correction on data read from the memory array and performs encryption and decryption of specific cipher algorithms, i.e., AES, SHA, and RSA, via a dedicated cryptographic engine. The SSD controller, comprising 1 to 4 flash memory controllers, is equipped with an embedded multicore processor. This processor operates the internal firmware known as the flash translation layer (FTL). The FTL is responsible for translating addresses between the logical memory space and the physical flash page, performing garbage collection (GC) to reclaim invalid or empty spaces, and executing wear leveling (WL) to ensure even memory usage across the memory blocks. The DRAM is utilized for storing a virtual-to-physical address mapping table and the firmware code of the FTL.
@@ II-C Threat Model
Modern commercial SSDs provide hardware-based data protection through SED feature, typically employing AES encryption with a dedicated engine in the SSD controller [68, 117]. These controllers also support cryptographic primitives such as SHA and RSA for integrity verification and firmware authentication [94, 70, 34, 102]. In current SSD architectures, encryption is typically performed in one of two ways: (i) by a hardware engine integrated into the SSD controller or (ii) in software on the host CPU when the required algorithm is not supported by the controller.
When the SSD controller performs encryption, the NAND flash stores encrypted data sent from the controller [47]. However, storing encrypted data in NAND does not eliminate all security risks. For example, remnants of encryption keys or plaintext data may persist in NAND flash due to WL, GC [121, 3], or unexpected power loss [2, 128]. These residual traces may be extracted through physical attacks using flash readers [68]. Moreover, off-the-shelf NAND chips lack built-in access control mechanisms, leaving them vulnerable to unauthorized access through malicious firmware [68].
When encryption is performed by the CPU, sensitive data can be exposed to adversaries through probing attacks on the data bus [122, 123]. Attackers may also analyze power consumption patterns [91, 90] or electromagnetic emissions [30, 97] using sensors or oscilloscopes to infer secret keys or other sensitive information. Furthermore, software-based encryption is at risk of microarchitectural attacks, such as cache timing side channels, which can leak sensitive intermediate results that depend on cryptographic secrets [124, 83].
@@ III-A1 Block Cipher Algorithm
Modern SSDs provide a strong security level by employing block cipher algorithms that enable fast encryption and decryption to protect data. Most commercial self-encrypting SSDs support fixed block cipher algorithms such as AES [42, 44, 52], following storage security standards such as FIPS 140-2/3 [73, 74], TCG Opal [118], and IEEE 1619 [42]. However, modern security requirements are becoming increasingly diverse. To protect data vital to national security and public welfare, and maintain cryptographic sovereignty, many countries have enacted stricter data protection laws. Accordingly, they mandate or recommend the use of national standard block cipher algorithms to protect sensitive data (e.g., personal identifiers, financial records, and military information) stored in storage media. For example, the United States, the European Union, South Korea, China and Russia require the use of AES/Triple DES [77], AES [28], ARIA/SEED/HIGHT [56], SM4 [33], and GOST [119], respectively. Japan and Turkey recommend Camellia [23] and KET [53] for securing public sector data. This growing diversity of security demands indicates that relying on a single fixed block cipher algorithm in SSDs is inadequate to address today’s heterogeneous requirements.
@@ III-A2 Public-Key Cryptographic Algorithm
Recent storage devices are equipped with advanced security features based on data integrity and digital signatures. Storage security standards such as TCG Opal [118] and NVMe-oF specify public-key-based authentication mechanisms as core components, leading modern SSDs to support a limited set of public-key cryptographic algorithms such as RSA and NIST ECC-crypto222ECC-crypto refers to Elliptic Curve Cryptography. (e.g., secp256r1) [94, 70, 34, 102]. However, these fixed-function designs inherently lack flexibility to accommodate diverse regional security requirements. For instance, Germany’s Federal Office for Information Security recommends ECC-crypto based on Brainpool curves [29], while China promotes its domestic SM2 algorithm in industry [105]. However, commercial SSDs generally do not support such region-specific cryptographic algorithms, limiting their applicability in regions with specific regulatory requirements.
Moreover, as storage devices are entrusted with preserving sensitive data over long periods, ensuring their long-term security against cryptographic threats is critical. From this perspective, the advent of quantum computing poses a threat to existing public-key cryptosystems, since quantum computers can solve the underlying mathematical problems such as integer factorization and discrete logarithms using Shor’s algorithm [101]. This raises the concern that today’s widely adopted public-key security mechanisms may become ineffective in the near future. To address this, NIST has led post-quantum standardization and selected Dilithium, FALCON, and SPHINCS+ as digital signature standards [75]. In response, supporting these PQC algorithms is emerging as a critical requirement for future storage systems that can withstand future cryptographic threats.
@@ III-B Limits of CPU-based Encryption
Modern SSDs have evolved beyond simple data storage devices into intelligent security platforms equipped with various built-in protection mechanisms. In addition to encrypting data to ensure confidentiality, these devices also support integrity-focused features such as secure boot and tamper-proof logging, which perform signature verification and detect unauthorized modifications during system initialization and log management. To implement these security functions effectively, different cryptographic algorithms may be selected depending on the application domain and security priorities. For example, SSDs deployed in security-critical sectors, such as defense, may employ robust signature schemes such as ECDSA [80] or upcoming post-quantum algorithms to ensure high assurance. On the other hand, SSDs in latency-critical environments like financial trading systems may use lightweight block ciphers such as HIGHT [37] or Camellia [4] to minimize encryption latency while still ensuring data protection.
However, off-the-shelf SSDs typically support only a limited set of algorithms at the hardware level [94, 70, 34, 102], requiring other cryptographic operations to be offloaded to the host processor [107]. This CPU-based processing often incurs significant latency, making it difficult to meet the stringent timing requirements of practical applications. To mitigate this issue, it is necessary to minimize reliance on host-side processing in order to reduce the latency overhead caused by offloading cryptographic operations to the host processor. FlashVault addresses this challenge by supporting in-NAND self-encryption, enabling high-performance execution of diverse cryptographic algorithms. This section motivates the need for in-NAND processing by analyzing the inefficiencies of CPU-based encryption in real-world scenarios.
@@ III-B1 Experimental Setup
To quantitatively evaluate the performance overhead of CPU-based encryption, we measured the latency of various cryptographic algorithms on a desktop system with an Intel Core i7-13700K (16 cores), 32 GB DDR5-5600 memory, and a 2TB SK Hynix Gold P31 M.2 NVMe SSD (PCIe 3.0 ×4), running Ubuntu 22.04 LTS. Block ciphers and public-key cryptographic algorithms were implemented using the Botan library [11], and PQC algorithms were evaluated using the openssl speed benchmark tool in conjunction with liboqs [81] and OpenSSL [82]. For block ciphers, we used fixed 4 KB input messages, whereas public-key and PQC algorithms were evaluated across a range of input sizes. In particular, the experiments on public-key algorithms were based on scenarios modeling the verification process of secure boot and the signature validation used in security logging systems such as tamper-proof logging. As part of the analysis of secure boot latency, we used the latency value reported in [116] for the boot code loading process. Latency was measured using the clock_gettime() function, with fsync() and posix_fadvise() used to reduce the effects of OS-level file caching. Each algorithm was evaluated 1,000 times per input size, and we report the average latency.

### v2

# Flexible In-NAND Cryptographic Processing for Secure Flash Storage

@@ I Introduction
Today, governments and private enterprises leverage personal data to deliver highly customized policies and services [oecd_data_utilization, mckinsey_data_collect, segment_data_collect]. In doing so, they often collect sensitive information closely tied to individual privacy such as financial records and biometric data. While such data serves as a critical resource for enabling personalized services, it also becomes a prime target for attacks, with severe consequences if leaked. For instance, in 2013 and 2017, massive data breaches at two major U.S. corporations, Target (a retailer) and Equifax (a credit reporting agency), respectively, exposed sensitive personal information of millions of individuals [equifax_leak, target_leak]. These incidents resulted in credit card fraud and identity theft, with financial damages reaching hundreds of billions of dollars. To mitigate such illegal use and data misuse, many countries including the United States, the European Union, Japan, and South Korea have enacted data protection laws that mandate encryption and appropriate technical safeguards when storing personal data on storage devices [ccpa, gdpr, japan_ppc, korea_act]. In parallel, the industry has adopted standards like TCG Opal [tcg_ssc] and IEEE 1667 [ieee_std1667] to enhance storage security.
@@
In accordance with data protection laws, many countries require or recommend that public institutions and private enterprises handling personal information adopt national standard cryptographic algorithms [ccpa, gdpr, japan_ppc, korea_act].
However, current commercial SSDs currently support only specific algorithms, e.g., AES, SHA, and RSA [samsung_secure_firmware, micron_secure_firmware, ibm_tls, SK_Hynix_ssd_product].
As a result, whenever unsupported algorithms are required, encryption is offloaded to the software stack on the host CPU [snia_cpu].
This software-based encryption incurs I/O latency and may increase system-wide response time [d_shield].
Such delays not only degrade the user experience in public and enterprise services but also cause significant disruption in latency-sensitive applications, such as high-frequency trading systems [hft_damage] and decentralized exchanges [blockchain_damage].
[Limitation 2] Unencrypted data may be stored in NAND flash memory. Modern SSDs are equipped with data encryption and SED features [firmware_attack, tcg_sed].
However, the activation of these features is typically determined by user configuration [dell_sed, qnap_sed].
@@
Even when encryption is enabled, residual traces of cryptographic keys may remain in NAND flash due to garbage collection (GC) and wear-leveling (WL) during the key revocation process [wear_leveling_attack1, adaptive_privacy_ssd].
In addition, if a power loss occurs, plaintext data residing in the DRAM of the SSD may be flushed to NAND flash [power_loss_date, power_loss_fast].
[Limitation 3] NAND flash lacks architectural protections against software-based and physical attacks. Since commercial SSDs do not support trusted execution environments within their NAND flash chips, attackers can manipulate the flash translation layer (FTL) mapping tables to distort the data organization or analyze FTL logs to retrieve or tamper with residual data that have been logically deleted but remain physically present [iceclave].
In addition, attackers can connect NAND readers to the flash chip to extract unencrypted data, or remnants of data left behind due to GC and WL [firmware_attack].
@@
For instance, plaintext data transmitted between the SSD and host may be intercepted through bus probing [anti_bus_probing, anti_bus_probing2], and cryptographic keys may be inferred or stolen through power analysis [power_attack, power_attack2] or electromagnetic analysis [em_attack, em_attack2] during CPU-based encryption processes.
@@
This paper focuses on the architecture of FlashVault and its performance. While security functionality is a fundamental requirement driven by legal and industrial standards, performance is also a critical factor for the practical deployment of secure storage systems in real-world environments [storage_energy_latency, iceclave, rif]. Moreover, the structural characteristics of FlashVault inherently mitigate security vulnerabilities observed in conventional systems (Section II-C).
@@ II-A NAND Flash Memory Technology
The memory cells in early-generation SSDs are based on floating gate transistors (FGTs). An FGT is a complementary metal-oxide-semiconductor (CMOS) technology capable of holding electrical charges in an isolated silicon conductive layer (Fig. 2-(a)) [overview_nvm_tech]. Early generation flash memory arrays have a 2D planar structure with FGTs placed side-by-side on the die (Fig. 2-(c)). However, as the process node has scaled down, FGTs encountered fabrication challenges, i.e., charge leakage caused by the thin conductive layer and intensified cell-to-cell interference [cell_interference, modern_ssd_survey]. To alleviate this limitation, FGTs were gradually superseded by transistors utilizing charge trap technology. These newer transistors employ a non-conductive silicon layer as a control gate, highly effective at retaining electrical charges (Fig. 2-(b)) [overview_nvm_tech, modern_ssd_survey]. Moreover, to meet demands for higher density, memory suppliers developed a 3D NAND structure (Fig. 2-(d)), featuring vertically stacked charge trap transistors alongside the peripheral circuit on the same integrated chip [samsung_3d_vnand, micron_3d_nand, Hynix_3d_vnand]. This 3D architecture significantly increased storage capacity by adding more layers achieving lower cost/bit. More recently, the industry has introduced 4D NAND flash products111In the industry, 4D V-NAND is also referred to as PUC (Periphery Under Cell) by SK hynix, CUA (CMOS Under Array) by Micron, and COP (Cell Over Periphery) by Samsung [tech_insight_survey]., which adopt an advanced vertical integration scheme to improve bit density (Fig. 2-(e)). These offer a reduced chip area compared to 3D NAND flash by positioning the peripheral circuit underneath the memory array [peri_area]. Fig. 3-(a) illustrates the area composition of 512 Gb-class 3D V-NAND chips from three major memory manufacturers, indicating that peripheral circuits account for approximately 2024% of the total chip area. By placing the peripheral circuits under the memory array, the 4D structure not only reduces the die footprint by a comparable amount (Fig. 3-(b)), but also leaves a portion of the bottom silicon layer unoccupied, since the relocated circuits are smaller in area than the memory array above. Motivated by this architectural characteristic, there have been efforts to leverage the resulting unused area by integrating auxiliary logic circuits, such as pattern matchers [pif], data processing units [4d_vnand_data_processing], and neuron-inspired circuits [4d_vnand_neuron_circuit]. This work leverages this area to embed reconfigurable cryptographic engines, enabling fast and secure near-data cryptographic processing.
@@ II-B Internal Architecture of SSDs
Fig. 4 illustrates the typical architecture of modern SSDs [modern_ssd_isca, modern_ssd_survey, modern_ssd_amber, decoupled_ssd]. An SSD consists of NAND flash packages, an SSD controller, and DRAM. Each NAND flash package contains multiple flash chips, each housing several memory planes. These planes are further divided into blocks and pages. A block is the smallest erasable unit, typically holding 1,024 pages. A page, the basic unit for read and program/write operations, usually has a size between 4 KB and 16 KB. Data read from or programmed to the flash memory is managed by a flash memory controller. This controller orchestrates data distribution to prevent bottlenecks among NAND packages sharing a byte-wide channel. It also performs error correction on data read from the memory array and performs encryption and decryption of specific cipher algorithms, i.e., AES, SHA, and RSA, via a dedicated cryptographic engine. The SSD controller, comprising 1 to 4 flash memory controllers, is equipped with an embedded multicore processor. This processor operates the internal firmware known as the flash translation layer (FTL). The FTL is responsible for translating addresses between the logical memory space and the physical flash page, performing garbage collection (GC) to reclaim invalid or empty spaces, and executing wear leveling (WL) to ensure even memory usage across the memory blocks. The DRAM is utilized for storing a virtual-to-physical address mapping table and the firmware code of the FTL.
@@ II-C Threat Model
Modern commercial SSDs provide hardware-based data protection through SED feature, typically employing AES encryption with a dedicated engine in the SSD controller [firmware_attack, tcg_sed]. These controllers also support cryptographic primitives such as SHA and RSA for integrity verification and firmware authentication [samsung_secure_firmware, micron_secure_firmware, ibm_tls, SK_Hynix_ssd_product]. In current SSD architectures, encryption is typically performed in one of two ways: (i) by a hardware engine integrated into the SSD controller or (ii) in software on the host CPU when the required algorithm is not supported by the controller.
When the SSD controller performs encryption, the NAND flash stores encrypted data sent from the controller [iceclave]. However, storing encrypted data in NAND does not eliminate all security risks. For example, remnants of encryption keys or plaintext data may persist in NAND flash due to WL, GC [wear_leveling_attack1, adaptive_privacy_ssd], or unexpected power loss [power_loss_date, power_loss_fast]. These residual traces may be extracted through physical attacks using flash readers [firmware_attack]. Moreover, off-the-shelf NAND chips lack built-in access control mechanisms, leaving them vulnerable to unauthorized access through malicious firmware [firmware_attack].
When encryption is performed by the CPU, sensitive data can be exposed to adversaries through probing attacks on the data bus [anti_bus_probing, anti_bus_probing2]. Attackers may also analyze power consumption patterns [power_attack, power_attack2] or electromagnetic emissions [em_attack, em_attack2] using sensors or oscilloscopes to infer secret keys or other sensitive information. Furthermore, software-based encryption is at risk of microarchitectural attacks, such as cache timing side channels, which can leak sensitive intermediate results that depend on cryptographic secrets [cache_attack1, cache_attack2].
@@ III-A1 Block Cipher Algorithm
Modern SSDs provide a strong security level by employing block cipher algorithms that enable fast encryption and decryption to protect data. Most commercial self-encrypting SSDs support fixed block cipher algorithms such as AES [ieee_aes_std, intel_aes, kingston_standard], following storage security standards such as FIPS 140-2/3 [fips_140_2, fips_140_3], TCG Opal [tcg_ssc], and IEEE 1619 [ieee_aes_std]. However, modern security requirements are becoming increasingly diverse. To protect data vital to national security and public welfare, and maintain cryptographic sovereignty, many countries have enacted stricter data protection laws. Accordingly, they mandate or recommend the use of national standard block cipher algorithms to protect sensitive data (e.g., personal identifiers, financial records, and military information) stored in storage media. For example, the United States, the European Union, South Korea, China and Russia require the use of AES/Triple DES [cmvp], AES [gdpr], ARIA/SEED/HIGHT [kcmvp_example], SM4 [sm4_report], and GOST [gost], respectively. Japan and Turkey recommend Camellia [japan_camellia] and KET [kvkk] for securing public sector data. This growing diversity of security demands indicates that relying on a single fixed block cipher algorithm in SSDs is inadequate to address today’s heterogeneous requirements.
@@ III-A2 Public-Key Cryptographic Algorithm
Recent storage devices are equipped with advanced security features based on data integrity and digital signatures. Storage security standards such as TCG Opal [tcg_ssc] and NVMe-oF specify public-key-based authentication mechanisms as core components, leading modern SSDs to support a limited set of public-key cryptographic algorithms such as RSA and NIST ECC-crypto222ECC-crypto refers to Elliptic Curve Cryptography. (e.g., secp256r1) [samsung_secure_firmware, micron_secure_firmware, ibm_tls, SK_Hynix_ssd_product]. However, these fixed-function designs inherently lack flexibility to accommodate diverse regional security requirements. For instance, Germany’s Federal Office for Information Security recommends ECC-crypto based on Brainpool curves [ecc_bsi], while China promotes its domestic SM2 algorithm in industry [china_sm2]. However, commercial SSDs generally do not support such region-specific cryptographic algorithms, limiting their applicability in regions with specific regulatory requirements.
Moreover, as storage devices are entrusted with preserving sensitive data over long periods, ensuring their long-term security against cryptographic threats is critical. From this perspective, the advent of quantum computing poses a threat to existing public-key cryptosystems, since quantum computers can solve the underlying mathematical problems such as integer factorization and discrete logarithms using Shor’s algorithm [quantum_threat]. This raises the concern that today’s widely adopted public-key security mechanisms may become ineffective in the near future. To address this, NIST has led post-quantum standardization and selected Dilithium, FALCON, and SPHINCS+ as digital signature standards [nist_pqc_altorithms]. In response, supporting these PQC algorithms is emerging as a critical requirement for future storage systems that can withstand future cryptographic threats.
@@ III-B Limits of CPU-based Encryption
Modern SSDs have evolved beyond simple data storage devices into intelligent security platforms equipped with various built-in protection mechanisms. In addition to encrypting data to ensure confidentiality, these devices also support integrity-focused features such as secure boot and tamper-proof logging, which perform signature verification and detect unauthorized modifications during system initialization and log management. To implement these security functions effectively, different cryptographic algorithms may be selected depending on the application domain and security priorities. For example, SSDs deployed in security-critical sectors, such as defense, may employ robust signature schemes such as ECDSA [ecdsa] or upcoming post-quantum algorithms to ensure high assurance. On the other hand, SSDs in latency-critical environments like financial trading systems may use lightweight block ciphers such as HIGHT [HIGHT] or Camellia [camellia] to minimize encryption latency while still ensuring data protection.
However, off-the-shelf SSDs typically support only a limited set of algorithms at the hardware level [samsung_secure_firmware, micron_secure_firmware, ibm_tls, SK_Hynix_ssd_product], requiring other cryptographic operations to be offloaded to the host processor [snia_cpu]. This CPU-based processing often incurs significant latency, making it difficult to meet the stringent timing requirements of practical applications. To mitigate this issue, it is necessary to minimize reliance on host-side processing in order to reduce the latency overhead caused by offloading cryptographic operations to the host processor. FlashVault addresses this challenge by supporting in-NAND self-encryption, enabling high-performance execution of diverse cryptographic algorithms. This section motivates the need for in-NAND processing by analyzing the inefficiencies of CPU-based encryption in real-world scenarios.
@@ III-B1 Experimental Setup
To quantitatively evaluate the performance overhead of CPU-based encryption, we measured the latency of various cryptographic algorithms on a desktop system with an Intel Core i7-13700K (16 cores), 32 GB DDR5-5600 memory, and a 2TB SK Hynix Gold P31 M.2 NVMe SSD (PCIe 3.0 ×4), running Ubuntu 22.04 LTS. Block ciphers and public-key cryptographic algorithms were implemented using the Botan library [botan], and PQC algorithms were evaluated using the openssl speed benchmark tool in conjunction with liboqs [liboqs] and OpenSSL [openssl]. For block ciphers, we used fixed 4 KB input messages, whereas public-key and PQC algorithms were evaluated across a range of input sizes. In particular, the experiments on public-key algorithms were based on scenarios modeling the verification process of secure boot and the signature validation used in security logging systems such as tamper-proof logging. As part of the analysis of secure boot latency, we used the latency value reported in [texas_boot_size] for the boot code loading process. Latency was measured using the clock_gettime() function, with fsync() and posix_fadvise() used to reduce the effects of OS-level file caching. Each algorithm was evaluated 1,000 times per input size, and we report the average latency.